OpenStudy (anonymous):

the cost of a family membership at a health club is $58 per month plus a one-time $129 start-up fee.if a family spent $651,how many months is their membership?

OpenStudy (anonymous):

58x+129=651 58x+129-129=651-129 58x=522 58/58x=522/58 x=9 So they had membership for 9 months

OpenStudy (anonymous):

Does that answer your question?

OpenStudy (anonymous):

yes thx